OpenStudy (anonymous):

cos(4x)-cos(6x)=0 find all values 0

OpenStudy (xapproachesinfinity):

you could use sum to product or what it is called

OpenStudy (xapproachesinfinity):

how did you achieve cos(4x)=4cosx there is not such thing

OpenStudy (anonymous):

I dont think you can do that @ageta

OpenStudy (xapproachesinfinity):

cos(4x)-cos(6x)=2sin(5x)sinx

OpenStudy (xapproachesinfinity):

set that equal to zero i used sum to product if you don't know i shall explain...

OpenStudy (xapproachesinfinity):

clear?

OpenStudy (anonymous):

Yes but this leaves me in a hole almost as large.. I am trying to get it down to a single sin(x) or cos(x) in order to find the angles that satisfy the question

OpenStudy (xapproachesinfinity):

what do you mean? you 2sin(5x)sinx=0 ===>sin5x=0 or sinx=0

OpenStudy (xapproachesinfinity):

yes?

OpenStudy (xapproachesinfinity):

it is not all the time a single sin or cos as you think

OpenStudy (xapproachesinfinity):

you could have tanxsin5xcos3x=0 this is very legitimate

OpenStudy (anonymous):

So how do I use sin(5x) to find the angles?

OpenStudy (xapproachesinfinity):

if you have product of terms equal to zero it means that each term can equal zero

OpenStudy (xapproachesinfinity):

like when you have (x+2)(x-3)(x-1)=0 =====>x=-2 or x=3 or x=1 the same thing applies to trig (it is just algebra again)

OpenStudy (xapproachesinfinity):

sin(5x)=0 you ask what angle in the range 0 to 2pi gives us 0 when applying sin to it that would be 0, pi, 2pi

OpenStudy (xapproachesinfinity):

yes?

OpenStudy (anonymous):

so the 5x is not significant?

OpenStudy (xapproachesinfinity):

5x is a whole quantity by it self it is acting exactly like x

OpenStudy (xapproachesinfinity):

it is not about significance! you have 5x you need to solve for x only so you got as i said 5x=0 or 5x=pi or 5x=2pi then solve for x

OpenStudy (xapproachesinfinity):

that's only for sin5x=0 after you move to sinx=0

OpenStudy (xapproachesinfinity):

eh i'm sorry 0 and 2pi are not included actually i didn't pay attention i thought it was 0<=x<=360

OpenStudy (xapproachesinfinity):

so it's only 5x=pi

OpenStudy (anonymous):

The answers for the angles that are results of cos(4x)-cos(6x)= 0 are: 0, pi/5, 2pi/5, 3pi/5, 4pi/5, pi, 6pi/5, 7pi/5, 8pi/5, 9pi/5,

OpenStudy (anonymous):

I meant 0<x<360 - -

OpenStudy (anonymous):

However i still do not see how one would arrive at all of those angles

OpenStudy (xapproachesinfinity):

don't worry about options you have 5x=pi ====>x=pi/5 that's one solution

OpenStudy (xapproachesinfinity):

hmm sum of them might not be answers since we are only taking from 0 to 360

OpenStudy (anonymous):

These are the answers in the textbook, I was just trying to find out how to arrive at them

OpenStudy (xapproachesinfinity):

for sin5x we found x=pi/5 for \(\large x\in(0,2\pi)\)

OpenStudy (anonymous):

I think the way to go is cos(2x+2x) -cos(4x+2x) and just to finish that out

OpenStudy (xapproachesinfinity):

now move on to sinx=0 ===> x=pi for \(\large x\in(0,2\pi)\)

OpenStudy (xapproachesinfinity):

to me it appear there only two solution to \(\large cos4x-cos6x=0\) when \(0<x<360 ~or~ 0<x<2\pi\)

OpenStudy (xapproachesinfinity):

unless if you did a mistake in typing the range of the solutions

OpenStudy (anonymous):

nope thats gread thank you very much!

OpenStudy (xapproachesinfinity):

No problem! never look at answer options they just confuse you

OpenStudy (xapproachesinfinity):

focus on solving first

OpenStudy (xapproachesinfinity):

and you need to get you sum to product formula memorized if you can't bring them back by any other technique they are very important

OpenStudy (xapproachesinfinity):

http://colalg.math.csusb.edu/~devel/IT/main/m05_identities/src/s04_sumtoproduct.html this might help